OpenStudy (anonymous):

Solve: 7a − 5 = 2(2a −13) A. 5/7 B. -31/3 C. -7 D. 7

OpenStudy (anonymous):

7a − 5 = 2(2a −13) 7a - 5 =4a-26 7a-4a=-26+5 3a=-21 a=-7

OpenStudy (anonymous):

7a-5=4a-26 7a-4a=5-26 >3a=-21 a=-7 So the answer is C :D

OpenStudy (amistre64):

it looks like you each did a good job, but they only have one medal to give out :/

OpenStudy (amistre64):

the key is, medals dont mean a thing :)

OpenStudy (anonymous):

@amistre64 :i'am not keen on getting medals ,i was just wonder what was ma answer missing ? :)

OpenStudy (anonymous):

jk ...

OpenStudy (amistre64):

i think its missing a cool virtual pet; like a ...whats those things called, digimon?

OpenStudy (anonymous):

i am not worry about medals XD medals is not important right?